The resistivity of a potentiometer wire is  , if the area of cross section of the wire is 4cm2 the current flowing in the circuit is 1A the potential gradient is 7.5 v/m

Grade:12

2 Answers

Khimraj
5 years ago
Potential gradient = p*I/A
So 7.5 = p*1/4*10-4
so p = 3*10-3 ohm*m
Hope it clears.....................
Shivangi Baral
5 years ago
let symbols have their usual meanings so,
V = i R = i pl/A
=> V/l = ip/A
=> 7.5 = 1 * p / ( 4 * 10^-4)
=> p = 3 * 10^-3 ohm m